# references.bib

### v2

@@ Abstract
Hierarchical reinforcement learning (HRL) is hypothesized to be able to take advantage of the inherent hierarchy in robot learning tasks with sparse reward schemes, in contrast to more traditional reinforcement learning algorithms. In this research, hierarchical reinforcement learning is evaluated and contrasted with standard reinforcement learning in complex navigation tasks. We evaluate unique characteristics of HRL, including their ability to create sub-goals and the termination function. We constructed experiments to test the differences between PPO and HRL, different ways of creating sub-goals, manual vs automatic sub-goal creation, and the effects of the frequency of termination on performance. These experiments highlight the advantages of HRL and how it achieves these advantages.
@@ 1 Introduction
Navigation is an essential part of robotics. Every robot needs to navigated whether it is an arm in a factory moving from part to part or a roomba navigating the room to clean. But navigation problems can be extremely complex, this research aims to find out how to deal with this complexity. The proposed solution to the navigation problem is reinforcement learning, but it can struggle with tasks that have sparse reward structures. One method to deal with this complexity is to build a hierarchy. This comes from nature as humans build hierarchies to solve problems. For example if someone was going to bake a cake, they would first break it into to parts, gathering the ingredients, mixing the ingredients, etc. This is a form of build a hierarchy by breaking the task down into smaller sub-tasks. This can be used in conjunction with reinforcement learning to better solve complex tasks.
Hierarchical reinforcement learning (HRL) is a form of reinforcement learning that aims to overcome some of the shortcomings of traditional RL in environments with sparse reward schemes by breaking down tasks into sub-tasks and creating a hierarchy. In this research, we study HRL’s ability to solve complex spatial navigation problems by forming sub-tasks. This research contributes to the scientific community by providing a comparison between standard reinforcement learning and hierarchical reinforcement learning, an evaluation of different ways to create sub-goals, a comparison of manual and automatic sub-goal creation, and a study of the frequency of sub-goals on performance.
@@ 2.1 Hierarchical Reinforcement Learning
There are many different types of reinforcement learning (RL) algorithms, including traditional “shallow” and hierarchical reinforcement learning (HRL). HRL is a type of reinforcement learning that breaks down tasks into sub-tasks, creating a hierarchy to solve a task. This design mimics how animals and humans break down tasks into smaller problems before completing a problem. For example, when baking a cake, a person would break it into smaller parts, like gathering the ingredients or mixing everything together. In HRL, this design is applied at solving tasks with sparse reward structures, which are typically more challenging for traditional reinforcement learning.
@@ 2.2 Types of Hierarchical Reinforcement Learning
There are multiple HRL algorithms, as summarized in Table 1 (see Shubham Pateria et al. [HRLSurvey] for a review).
@@
Our research is focused on single-agent, single-task problems with sub-task discovery. In this category, there are two main areas of research: the option architecture and feudal reinforcement learning [SUTTON] [FRL]. A third area is the MAXQ method [MAXQ], which has lost popularity over time as it has too many requirements.
@@ 2.3 Options
One of the most popular types of Hierarchical Reinforcement Learning is the Option Architecture developed by Sutton et al. [SUTTON]. The Option Architecture is a way to introduce temporal abstraction to a Markov Decision Process (MDP). Introducing time violates the assumptions of an MDP, making such a process a Semi-Markov Decision Process (SMDP). SMDPs are difficult to solve. For this purpose, Sutton et al. [SUTTON] introduced the Option Architecture, which allows for an SMDP to be approximated as an MDP while still incorporating temporal abstraction. This is accomplished by replacing actions in the MDP with options, where an option is an action temporally extended over several states, as illustrated in Figure 1. An option is comprised of three parts: the initiation set, the policy, and the termination set. The initiation set is a subset of the state space where the option can be selected and used. The policy is the actions taken during the selected option. The termination set is a subset of the state space where the option will end. An example of how this would look in robotics is if a robot vacuum is cleaning a house with three rooms, then each room would be an option, with the termination set being the door to the next room. The initiation set would be the door to the current room. The option policy would be the policy to clean the current room. Accordingly, each option is an independent MDP, with a policy over the options that is approximated as an MDP with the actions being options.
@@ 2.3.1 Option Critic Architecture
The Option Critic Architecture builds upon the Option Architecture [SUTTON] by incorporating an algorithm to create the options using equations based on the Policy Gradient Methods [PolicyGradient]. To simplify the creation of the options, the architecture sets the initiation set to all states, so the initiation set does not need to be learned [OptionCritic]. This has several advantages and disadvantages, the primary advantage being it simplifies the algorithm as it does not need to learn the initiation set, with the primary disadvantage being it removes some of the functionality of the Option Architecture, making the role of the policy over options more difficult as it has to evaluate every option in every state.
@@ 2.3.2 Options of Interest
The Interest Option Critic (IOC) Architecture is another algorithm built on top of the Option Critic Architecture [OptionCritic]. When the Option Critic Architecture simplifies the problem by setting the initiation set to all possible states. The Interest Option Critic instead replaces the initiation set with a separate interest function for each option [IOC]. This reduces the complexity but maintains the advantages of having an initiation set. Using the interest function makes the options specialized, resulting in faster learning and better adaptation [IOC].
@@ 2.3.3 Attention Option Critic
Attention Option Critic (AOC) is another algorithm that builds on top of the Option Critic Architecture [OptionCritic] [AOC]. It is similar to IOC [IOC], in that it adds some form of initiation set to the Option Critic Architecture [OptionCritic]. In this case, instead of using the interest function, the AOC uses an attention mechanism to act as an initiation set and to determine when each option is used [attenpaper]. This is done to make the options diverse, stable, and transferable [AOC].
@@ 2.3.4 Deep Skill Chaining
Although the Option Critic Architecture [OptionCritic] is the predominant option architecture in the field of hierarchical reinforcement learning, another approach to using options is Deep Skill Chaining [DSC]. Deep Skill Chaining is a form of HRL that chains options together to solve tasks. Instead of initiation at the start state, DSC chains the skills backwards starting from the goal until it reaches the original state. In contrast to OC [OptionCritic], DSC does learn an initiation set for each option in the algorithm. This can have advantages such as making it so the option does not have to learn a policy for every state in the state space. In contrast, some disadvantages exist, such as making the algorithm more complex [DSC].
@@ 2.4 Feudal Algorithms
Another type of Hierarchical Reinforcement Learning is Feudal Reinforcement Learning [FRL]. The goal of Feudal Reinforcement Learning is to speed up learning by learning at multiple resolutions in space and time [FRL]. It does this by having multiple levels and learning at each one. It starts by creating a hierarchy with a high level manager who manages lower level sub-managers who, in turn, manage their own lower level managers. This goes on for as many layers as decided by the designer. The lowest level manager operates on a portion of the state space and outputs to action to the environment. The manager on top of the lowest level operates on an abstract state space and as levels increase in the state becomes more abstract. The higher level managers pick the lower level managers based on the their state abstractions. This structure can speed up learning and lead to better performance model.
@@ 2.4.1 Feudal Networks for Hierarchical Reinforcement Learning
The FeUdal Networks for Hierarchical Reinforcement Learning (FUN) is an algorithm built on top of the Feudal Algorithms [FRL] [FUN]. In this algorithm, instead of having a variable number of layers as in Feudal Reinforcement Learning, the FUN algorithm only has two levels. This is done to simplify the algorithm and make it have the same number of layers for any problem. In this algorithm, the top layer is known as the manager, and the lower level is known as the worker [FUN]. The worker chooses the primitive actions in the environment and accesses the environment’s state space directly. The manager works in an abstract state space and sends goals to the worker. The worker then uses the primitive actions to achieve the goal given by the manager. This allows the worker to work with short term goals, while the manager can focus on long term goals. This structure makes FUN better on long-term credit assignments than traditional algorithms.
@@ 2.4.2 Hierarchical Reinforcement Learning With Off-Policy Correction
HIerarchical Reinforcement learning with Off-policy correction (HIRO) is another algorithm with two levels of controllers [HIRO]. The goal of HIRO is to be a general off-policy hierarchical reinforcement learning algorithm. This algorithm is similar to FUN [FUN], but one of the major differences is that HIRO does not abstract the state when communicating between the two levels [HIRO]. The other major difference is that HIRO is an off-policy algorithm while FUN is on-policy.
@@ 2.5 MAXQ
The MAXQ Decomposition algorithm was a popular algorithm around the time it came out in the 1990s [MAXQ]. The MAXQ algorithm requires the designer to define sub-goals and sub-tasks to complete these goals. The MAXQ then makes the hierarchy and uses it to decompose the value function. The value function is then used to create the policy for the task.
@@ 3.1 Policy Gradient Methods
Policy Gradient Methods [PolicyGradient] are a popular family of reinforcement learning algorithms that derive equations to represent the policy with a function approximator. This algorithm is one of the first reinforcement learning algorithms and is the foundation for many state-of-the-art algorithms. The update function derived for this algorithm is shown in equation 1 and is the basis of both the Option Critic Architecture [OptionCritic] and Proximal Policy Optimization [PPO].
@@ 3.2 Options
The options model used for this research is based on Options from Sutton, Precup, and Singh [SUTTON]. In their work the authors proposed using options in reinforcement learning to build a hierarchy and introduce temporal abstraction into the algorithm. An option is based on the idea of an action over time; i.e., a temporally extended action. A policy can then use these options as actions and operate on a higher temporal plane while still acting like an MDP. Each option $\omega$, is a set of $(I_{\omega},\pi_{\omega},\beta_{\omega})$. The $I_{\omega}$ is the initial set that determines when the option starts. The $\pi_{\omega}$ is the policy for the option, and the $\beta_{\omega}$ is the termination set determining when the option should end.
@@ 3.3 Option Critic Architecture
The model used in these experiments is the Option Critic Architecture [OptionCritic]. This architecture is built around options from Sutton et al. [SUTTON]. The Option Critic Architecture has a select number of options, a policy over options, and a critic $Q_{U}$ as shown in Figure 2.
@@
The Option Critic Architecture itself provides no guarantee of convergence, but it is built off of the Option Architecture, which guarantees the convergence of the option policy if every action in every state is executed infinitely often [SUTTON]. There is one main reason this guarantee does not carry over into the Option Critic Architecture, which is that the Option Architecture does not learn the sub-goals and assumes they are set beforehand. Since the Option Critic Architecture learns the sub-goals, the guarantee of convergence does not apply. Meaning if the sub-goals are not correct, the algorithm may not converge, but in practice, this is rare as the algorithm is fairly robust.
@@ 3.4 Option Critic Hyperparameters
The hyperparameters used for these experiments are taken directly from Bacon, Harb, and Precup [OptionCritic]. These parameters are shown in Table 2.
@@
  • Freeze Interval is how often the second critic is updated as two critics were implemented to prevent critic overestimation from Hasselt [DoubleQ].
@@ 3.6 Proximal Policy Optimization
The Proximal Policy Optimization (PPO) algorithm [PPO], is a popular standard reinforcement learning algorithm. This algorithm is used as the baseline standard reinforcement learning algorithm to compare with hierarchical reinforcement learning. PPO is based on the policy gradient methods [PolicyGradient], but is optimized to do multiple mini-batch updates instead of one update per sample. Another important part of PPO is the clipped surrogate loss in Equation 6, which prevents the algorithm from learning more than $1+\epsilon$ in one step. To prevent it from taking too large of steps during convergence. This makes it an excellent algorithm in general and a perfect candidate to be a baseline as a state-of-the-art standard reinforcement learning algorithm.
@@
The neural network used for PPO in this research is shown in Figure 5. It has four fully connected layers will a ReLU in between each one.

### v3

# Hierarchical Reinforcement Learning in Multi-Goal Spatial Navigation with Autonomous Mobile Robots

@@ Abstract
Hierarchical reinforcement learning (HRL) is hypothesized to be able to leverage the inherent hierarchy in learning tasks where traditional reinforcement learning (RL) often fails. In this research, HRL is evaluated and contrasted with traditional RL in complex robotic navigation tasks. We evaluate unique characteristics of HRL, including its ability to create sub-goals and the termination functions. We constructed a number of experiments to test: 1) the differences between RL proximal policy optimization (PPO) and HRL, 2) different ways of creating sub-goals in HRL, 3) manual vs automatic sub-goal creation in HRL, and 4) the effects of the frequency of termination on performance in HRL. These experiments highlight the advantages of HRL over RL and how it achieves these advantages.
keywords:
Hierarchical Reinforcement Learning , Robotics , Navigation , Options Architecture , Machine Learning , Artificial Intelligence
journal: Nuclear Physics B
[…]
organization=University of South Florida,addressline=4202 E. Fowler Avenue, city=Tampa, postcode=33620, state=Florida, country=United States of America
{highlights}
A comparison between standard reinforcement learning and hierarchical reinforcement learning.
An evaluation of different ways to create sub-goals.
A comparison between manual and automatic sub-goal creation.
A study of the frequency of sub-goals on performance.
@@ 1 Introduction
Navigation is an essential part of robotics. Navigation problems can be extremely complex. This research aims to find out how to deal with this complexity. The proposed solution to the navigation problem involves reinforcement learning (RL), but it can struggle with tasks that have sparse reward structures dqn_human (26) hindsight_rl (32) reverse_rl (12) hindsight_replay (2). Traditional RL is unreliable and often fails when solving tasks with long-term reward schemes. If the rewards in a task are few with many actions in between, as in a long-term reward scheme, RL will often be unreliable when temporal difference is used to train the algorithm. Temporal difference can fail in these tasks as the reward will decay, causing the temporal difference equations to fail because the reward signal becomes too small. Additionally, when using neural networks as part of the algorithm the weakened reward signal can cause decaying gradients, preventing sufficient learning. Another problem for RL in sparse reward tasks is that the algorithm needs to complete a complex sequence of actions before reaching a non-zero reward. This causes the agent to rarely receive positive rewards, as random exploration will rarely complete the sequence correctly hindsight_rl (32).
One method to address this complexity is to build a hierarchy. This is inspired by nature as humans build hierarchies to solve problems. For example, if baking a cake, a person would first break it into parts: gathering the ingredients, mixing the ingredients, etc. This is a form of building a hierarchy by breaking down the task into smaller sub-tasks. This can be used in conjunction with RL to better solve complex tasks.
Hierarchical reinforcement learning (HRL) is a form of reinforcement learning that aims to overcome some of the shortcomings of traditional RL in environments with sparse reward schemes by breaking down tasks into sub-tasks and creating a hierarchy. Creating this hierarchy can lead to several advantages over traditional RL, including faster training, improved adaptability, and the ability to solve sparse reward problems.
In this research, we investigate HRL’s ability to solve complex robot spatial navigation problems by creating sub-tasks while analyzing different sub-goal creation strategies. The paper provides the following analysis: 1) a comparison between traditional RL and HRL, 2) an evaluation of different methods for creating sub-goals, 3) a comparison of manual and automatic sub-goal creation, and 4) a study of the frequency of sub-goals on performance.
Each experiment is designed to highlight a different aspect of hierarchical reinforcement learning, thereby illustrating key components of HRL:
  • 1.
    Experiment 1 is designed to show how HRL can solve sparse reward tasks when traditional RL fails. This is important because sparse reward tasks can be particularly difficult for RL, but many robotics tasks are sparse, so it is an important issue to study.
  • 2.
    Experiment 2 is designed to test different ways to create sub-goals and their importance in HRL. Although HRL was designed to build hierarchies to outperform standard reinforcement learning, some studies suggest this is not the reason for its improved performance HRLSometimes (28). This experiment is designed to test if this assumption is correct. But to fully test this, more than one experiment was required. Experiment four is also designed to test this assumption.
  • 3.
    Experiment 3 is designed to evaluate manual versus automatic sub-goals creation. In most cases, manual sub-goals may be infeasible, but in certain cases, they may be practical and needed if they improve performance. The third experiment will assess whether they improve the performance of the agent.
  • 4.
    Experiment 4 is designed to test how the frequency of sub-goals impacts performance. This is important as it shows the optimal number of sub-goals. Additionally, in conjunction with experiment two, it demonstrates the importance of sub-goals in HRL algorithms. These experiments show this by evaluating the result of removing sub-goals from the algorithm. This is important as it explains why HRL is able to outperform traditional reinforcement learning in sparse reward tasks.
@@ 2.1 Hierarchical Reinforcement Learning
There are many different types of reinforcement learning (RL) algorithms, including traditional “shallow” and hierarchical reinforcement learning (HRL). HRL is a type of reinforcement learning that breaks down tasks into sub-tasks, creating a hierarchy to solve a task. This design mimics how animals and humans break down tasks into smaller problems before completing a problem. In HRL, this design is applied to solving tasks with sparse reward structures, which are typically more challenging for traditional reinforcement learning algorithms dqn_human (26) hindsight_rl (32) reverse_rl (12) hindsight_replay (2).
@@ 2.2 Types of Hierarchical Reinforcement Learning
There are multiple HRL algorithms, as summarized in Table 1 (see Shubham Pateria et al. HRLSurvey (31) for a review).
@@
Our research is focused on single-agent, single-task problems with sub-task discovery. In this category, there are two main areas of research: the option architecture and feudal reinforcement learning SUTTON (40) FRL (9). A third area is the MAXQ method MAXQ (10), which has lost popularity over time as it has too many requirements.
@@ 2.3 Options
One of the most popular types of Hierarchical Reinforcement Learning is the Option Architecture developed by Sutton et al. SUTTON (40). The Option Architecture is a way to introduce temporal abstraction to a Markov Decision Process (MDP) sutton_rl_intro (39). Introducing time violates the assumptions of an MDP, making such a process a Semi-Markov Decision Process (SMDP). SMDPs are difficult to solve. For this purpose, Sutton et al. SUTTON (40) introduced the Option Architecture, which allows for an SMDP to be approximated as an MDP while still incorporating temporal abstraction. This is accomplished by replacing actions in the MDP with options, where an option is an action temporally extended over several states, as illustrated in Figure 1. An option is comprised of three parts: the initiation set, the policy, and the termination set. The initiation set is a subset of the state space where the option can be selected and used. The policy is the actions taken during the selected option. The termination set is a subset of the state space where the option will end. An example of how this would look in robotics is if a robot vacuum is cleaning a house with three rooms, then each room would be an option, with the termination set being the door to the next room. The initiation set would be the door to the current room. The option policy would be the policy to clean the current room. Accordingly, each option is an independent MDP, with a policy over the options that is approximated as an MDP with the actions being options. The Option Architecture does not learn the sub-goals and assumes they are preset but there are many algorithms that build on top of the Option Architecture to solve this problem OptionCritic (3) TerminationCritic (14) DSC (4) automatic_discovery (22) qcut (24).
@@ 2.3.1 Option Critic Architecture
The Option Critic Architecture builds upon the Option Architecture SUTTON (40) by incorporating an algorithm to create the options using equations based on the Policy Gradient Methods PolicyGradient (41). It also uses an actor critic structure from the Actor-Critic Algorithms ACA (18) the options being the actor and critic used to evaluate them. To simplify the creation of the options, the architecture sets the initiation set to all states, so the initiation set does not need to be learned OptionCritic (3). This has several advantages and disadvantages, the primary advantage being it simplifies the algorithm as it does not need to learn the initiation set, with the primary disadvantage being it removes some of the functionality of the Option Architecture, making the role of the policy over options more difficult as it has to evaluate every option in every state. There are several algorithms that are built on top of the Option Critic Architecture PPOOptionCritic (17) DSC (4), including some discussed in future sections IOC (16) AOC (8).
@@ 2.3.2 Options of Interest
The Interest Option Critic (IOC) Architecture is another algorithm built on top of the Option Critic Architecture OptionCritic (3). When the Option Critic Architecture simplifies the problem by setting the initiation set to all possible states. The Interest Option Critic instead replaces the initiation set with a separate interest function for each option IOC (16). This reduces the complexity but maintains the advantages of having an initiation set. Using the interest function makes the options specialized, resulting in faster learning and better adaptation IOC (16).
@@ 2.3.3 Attention Option Critic
Attention Option Critic (AOC) is another algorithm that builds on top of the Option Critic Architecture OptionCritic (3) AOC (8). It is similar to IOC IOC (16), in that it adds some form of initiation set to the Option Critic Architecture OptionCritic (3). In this case, instead of using the interest function, the AOC uses an attention mechanism to act as an initiation set and to determine when each option is used attenpaper (5). This is done to make the options diverse, stable, and transferable AOC (8).
@@ 2.3.4 Deep Skill Chaining
Although the Option Critic Architecture OptionCritic (3) is the predominant option architecture in the field of hierarchical reinforcement learning, another approach to using options is Deep Skill Chaining DSC (4). Deep Skill Chaining is a form of HRL that chains options together to solve tasks and is based on Skill Chaining Skill (19). Instead of initiation at the start state, DSC chains the skills backwards starting from the goal until it reaches the original state. In contrast to OC OptionCritic (3), DSC does learn an initiation set for each option in the algorithm. This can have advantages such as making it so the option does not have to learn a policy for every state in the state space. In contrast, some disadvantages exist, such as making the algorithm more complex DSC (4).
@@ 2.4 Feudal Algorithms
Another type of Hierarchical Reinforcement Learning is Feudal Reinforcement Learning FRL (9). The goal of Feudal Reinforcement Learning is to speed up learning by learning at multiple resolutions in space and time FRL (9). It does this by having multiple levels and learning at each one. It starts by creating a hierarchy with a high level manager who manages lower level sub-managers who, in turn, manage their own lower level managers. This goes on for as many layers as decided by the designer. The lowest level manager operates on a portion of the state space and outputs to action to the environment. The manager on top of the lowest level operates on an abstract state space and as levels increase in the state becomes more abstract. The higher level managers pick the lower level managers based on the their state abstractions. This structure can speed up learning and lead to better performance model.
@@ 2.4.1 Feudal Networks for Hierarchical Reinforcement Learning
The FeUdal Networks for Hierarchical Reinforcement Learning (FUN) is an algorithm built on top of the Feudal Algorithms FRL (9) FUN (42). In this algorithm, instead of having a variable number of layers as in Feudal Reinforcement Learning, the FUN algorithm only has two levels. This is done to simplify the algorithm and make it have the same number of layers for any problem. In this algorithm, the top layer is known as the manager, and the lower level is known as the worker FUN (42). The worker chooses the primitive actions in the environment and accesses the environment’s state space directly. The manager works in an abstract state space and sends goals to the worker. The worker then uses the primitive actions to achieve the goal given by the manager. This allows the worker to work with short term goals, while the manager can focus on long term goals. This structure makes FUN better on long-term credit assignments than traditional algorithms.
@@ 2.4.2 Hierarchical Reinforcement Learning With Off-Policy Correction
HIerarchical Reinforcement learning with Off-policy correction (HIRO) is another algorithm with two levels of controllers HIRO (27). The goal of HIRO is to be a general off-policy hierarchical reinforcement learning algorithm. This algorithm is similar to FUN FUN (42), but one of the major differences is that HIRO does not abstract the state when communicating between the two levels HIRO (27). The other major difference is that HIRO is an off-policy algorithm while FUN is on-policy.
@@ 2.5 MAXQ
The MAXQ Decomposition algorithm was a popular algorithm around the time it came out in the 1990s MAXQ (10). The MAXQ algorithm requires the designer to define sub-goals and sub-tasks to complete these goals. The MAXQ then makes the hierarchy and uses it to decompose the value function. The value function is then used to create the policy for the task.
@@ 3.1 Policy Gradient Methods
Policy Gradient Methods PolicyGradient (41) are a popular family of reinforcement learning algorithms that derive equations to represent the policy with a function approximator. This algorithm is one of the first reinforcement learning algorithms and is the foundation for many state-of-the-art algorithms a3c (25) ddpg (21) pgq (30) mpo (1) sac (13). The update function derived for this algorithm is shown in equation 1 and is the basis of both the Option Critic Architecture OptionCritic (3) and Proximal Policy Optimization PPO (34).
@@ 3.2 Options
The options model used for this research is based on Options from Sutton, Precup, and Singh SUTTON (40). In their work the authors proposed using options in reinforcement learning to build a hierarchy and introduce temporal abstraction into the algorithm. An option is based on the idea of an action over time; i.e., a temporally extended action. A policy can then use these options as actions and operate on a higher temporal plane while still acting like an MDP. Each option $\omega$, is a set of $(I_{\omega},\pi_{\omega},\beta_{\omega})$. The $I_{\omega}$ is the initial set that determines when the option starts. The $\pi_{\omega}$ is the policy for the option, and the $\beta_{\omega}$ is the termination set determining when the option should end.
@@ 3.3 Option Critic Architecture
The model used in these experiments is the Option Critic Architecture OptionCritic (3). This architecture is built around options from Sutton et al. SUTTON (40). The Option Critic Architecture has a select number of options, a policy over options, and a critic $Q_{U}$ as shown in Figure 2.
@@
The Option Critic Architecture itself provides no guarantee of convergence, but it is built off of the Option Architecture, which guarantees the convergence of the option policy if every action in every state is executed infinitely often SUTTON (40). There is one main reason this guarantee does not carry over into the Option Critic Architecture, which is that the Option Architecture does not learn the sub-goals and assumes they are set beforehand. Since the Option Critic Architecture learns the sub-goals, the guarantee of convergence does not apply. Meaning if the sub-goals are not correct, the algorithm may not converge, but in practice, this is rare as the algorithm is fairly robust.
@@ 3.4 Option Critic Hyperparameters
The hyperparameters used for these experiments are taken directly from Bacon, Harb, and Precup OptionCritic (3). These parameters are shown in Table 2.
@@
  • 7.
    Freeze Interval is how often the second critic is updated as two critics were implemented to prevent critic overestimation from Hasselt DoubleQ (15).
  • 8.
@@ 3.6 Proximal Policy Optimization
The Proximal Policy Optimization (PPO) algorithm PPO (34), is a popular standard reinforcement learning algorithm. This algorithm is used as the baseline standard reinforcement learning algorithm to compare with hierarchical reinforcement learning. PPO is based on the policy gradient methods PolicyGradient (41) and trust region policy optimization trpo (35), but is optimized to do multiple mini-batch updates instead of one update per sample. Another important part of PPO is the clipped surrogate loss in Equation 6, which prevents the algorithm from learning more than $1+\epsilon$ in one step. To prevent it from taking too large of steps during convergence. This makes it an excellent algorithm in general and a perfect candidate to be a baseline as a state-of-the-art standard reinforcement learning algorithm ppo_a3c (33) ppo_trpo (11) ppo_car (38) ppo_carv2 (37) ppo_uav (20).
@@
The neural network used for PPO in this research is shown in Figure 5. It has four fully connected layers with a ReLU in between each one relu_layer (29).